3. Show that the mass of radium (Mass number= 226, half-life = 1600 yrs) with an


activity of a curie (1Ci) is almost 1gm.



1
Expert's answer
2021-12-27T13:03:58-0500

"A=\frac m M N_A\frac{\ln 2}{T^{1/2}},\implies"

"m=\frac{AMT^{1/2}}{N_A \ln2},"

"m=\frac{1\cdot 3.7\cdot 10^{10}\cdot226\cdot 1600\cdot 365\cdot 86400}{6.02\cdot 10^{23}\cdot0.69}=1~g."